# Automatic Classification of Subjective Time Perception Using Multi-modal Physiological Data of Air Traffic Controllers

### v1

@@ Abstract
One indicator of well-being can be the person’s subjective time perception. In our project ChronoPilot, we aim to develop a device that modulates human subjective time perception. In this study, we present a method to automatically assess the subjective time perception of air traffic controllers, a group often faced with demanding conditions, using their physiological data and eleven state-of-the-art machine learning classifiers. The physiological data consist of photoplethysmogram, electrodermal activity, and temperature data. We find that the support vector classifier works best with an accuracy of 79 % and electrodermal activity provides the most descriptive biomarker. These findings are an important step towards closing the feedback loop of our ChronoPilot-device to automatically modulate the user’s subjective time perception. This technological advancement may promise improvements in task management, stress reduction, and overall productivity in high-stakes professions.
@@ I Introduction
Figure 1: Fundamental control loop of the envisioned ChronoPilot-device. TP Input is the user’s desired time perception for maximal productivity and well-being. The Controller produces Stimuli and applies them to the user. During the stimulation, physiological data are collected, which are used as feedback for the ChronoPilot-device to adjust the stimuli accordingly. Our objective is to develop machine learning methods, which are able to provide this feedback by classifying the perceived passage of time based on physiological data.
@@
The only study we are aware of classifying subjective perceived passage of time based on physiological signals was done by Orlandic et al. [11]. They stimulated users with different video clips selected from the Emotional Movie [14] and FilmStim [15] databases, which lead to different emotional states (e.g., neutral, fear, sadness) along with cognitive tasks, such as mathematical tasks or the Stroop Color Test (varies cognitive load through congruent and incongruent stimuli). They collected skin temperature, electrodermal activity, respiration, electrocardiogram, and photoplethysmogram data. The feature extraction was done on non-overlapping 45-second slices, which were normalized by subtracting the mean and dividing by the standard deviation of the training dataset. They then applied a set of machine learning classification algorithms. We go beyond their approach as we use data of a more complex real-world scenario, namely flight control, instead of an in-lab experimental setup. We assess the perceived passage of time using a questionnaire instead of the relative estimation error. Our two main contributions are: (1) an automatic multi-modal subjective perceived passage of time perception estimation technique, using photoplethysmogram, electrodermal activity, and temperature data measured by of-the-shelf wearable sensors and (2) a comparison between eleven state-of-the-art machine learning classifiers for subjective perceived passage of time classification in combination with feature selection methods and feature importance analysis using the SHapley Additive exPlanations (SHAP) [16] framework. We show generalizability by validating our methods using leave-one-subject-out cross validation (LOSOCV). We obtain accuracies of up to 79 % on real-world physiological data obtained during flight controller helicopter training sequences.
@@ 1 H, GR1 H, EN2 H, GR2 H, ENvery slowslowmediumfastvery fastslowfast
Figure 2: Violin plots of the perceived passage of time labels of the different settings of the flight controller experiment (H: helicopter number, GR: Greek, EN: English; $n=12$). The left side of the violin plot (orange) displays the scaled label distribution from the questionnaire (scale is the left axis). The right side (blue) shows the label distribution after thresholding (dashed line) in the slow and fast class. The dots represent individual samples. We observe that depending on the number of helicopters the majority of the controllers perceived the passage of time as slow when there is a one helicopter and fast when there are two helicopters.
For classification, we use the perceived passage of time estimation from the controllers’ questionnaires obtained at the end of the experiment. To allow for better comparability between questionnaire answers, we min-max-scale each controller’s answers from “very slow” to “very fast”, see Fig. 2, left half (orange). We sort the settings by speaking in Greek and one helicopter to speaking in English and two helicopters as we suspect an increase in mental load for speaking in a non-native language and an increased number of to-be-controlled objects. This sorting suggest a binary classification problem, namely, “slow” for one helicopter settings and “fast” for two helicopters settings. By thresholding the scaled labels at medium ($>3$ is labeled as fast), we obtain a binary classification problem with 26 sequences of class slow and 22 sequences of class fast, see Fig. 2, right half (blue).
@@ IV-4 Limitations
Considering the future application scenario of providing feedback to a ChronoPilot-device, we encounter several practical limitations for data collection. Depending on the actual use case, users may be required to be mobile, see for example, the industrial production scenario [3]. Motion artifacts may corrupt the physiological data, such as PPG. Hence, we may limit ourselves to physiological data that can be collected even if users are mobile and in motion. One possible PPG replacement is measuring electrocardiography (ECG), which directly measures the heart electrical activity. ECG allows for extracting the same features as PPG. In addition, we could derive respiration features from ECG (also called EDR), which is another often used non-invasively measurable physiological signal. Using a thermopile as sensor also highly depends on the later use case, for example, considering again the industrial production scenario the user has to stand or walk around. One possible replacement could be measuring the skin and surrounding temperature. Skin temperature can often be measured together with EDA using the same device. Our results indicate that EDA is a strong descriptive feature and should definitely be used in further studies. Another interesting information source could be eye tracking, as it provides information about the user’s focus [24] and it could be integrated in a wearable ChronoPilot-device, such as VR/AR headsets.
@@ V Conclusion
In the present study, we were able to show that it is possible to estimate the perceived passage of time of air traffic controllers based on physiological data, such as PPG, EDA, and temperature measurements of a thermopile during helicopter training sequences of varying workload conditions with high accuracy up to 79 %. EDA seems to provide the most descriptive features based on SHAP analysis. This finding reinforces our assumption that we can influence the perceived passage of time by varying mental workload given that EDA is considered a biomarker for mental workload and stress. In terms of our objective of building a ChronoPilot-device, we have shown that we can use physiological data as feedback to close the control loop. For our future work, larger datasets may improve the likelihood of generalizability. This can be done in two ways: increasing the number of participants or the amount of data per participant. In addition, larger datasets would allow us to explore deep learning methods. The literature on exploiting physiological data to automatically detect a person’s subjective time perception is scarce. To date, a closed-loop approach for guided manipulation of time perception has not yet been attempted. Our results reported here encourage us in our effort of creating an automatic closed-loop approach for modulating subjective time perception in real-world scenarios as envisioned in the ChronoPilot project. By exploiting the plasticity of subjective time perception, we can develop future technology that enhances productivity and well-being.

### v2

@@ Abstract
In high-pressure environments where human individuals must simultaneously monitor multiple entities, communicate effectively, and maintain intense focus, the perception of time becomes a critical factor influencing performance and well-being. One indicator of well-being can be the person’s subjective time perception. In our project ChronoPilot, we aim to develop a device that modulates human subjective time perception. In this study, we present a method to automatically assess the subjective time perception of air traffic controllers, a group often faced with demanding conditions, using their physiological data and eleven state-of-the-art machine learning classifiers. The physiological data consist of photoplethysmogram, electrodermal activity, and temperature data. We find that the support vector classifier works best with an accuracy of 79 % and electrodermal activity provides the most descriptive biomarker. These findings are an important step towards closing the feedback loop of our ChronoPilot-device to automatically modulate the user’s subjective time perception. This technological advancement may promise improvements in task management, stress reduction, and overall productivity in high-stakes professions.
@@ I Introduction
Figure 1: Fundamental control loop of the envisioned ChronoPilot-device. TP Input is the user’s desired time perception for maximal productivity and well-being. The Controller produces Stimuli and applies them to the user. Physiological data are collected as feedback for the ChronoPilot-device to adjust the stimuli accordingly. Our objective is to develop machine learning methods, which are able to provide this feedback by classifying the perceived passage of time based on physiological data.
@@
The only study we are aware of that is classifying subjective time perception based on physiological signals was done by Orlandic et al. [11]. They stimulated users with different video clips selected from the Emotional Movie [14] and FilmStim [15] databases, which lead to different emotional states (e.g., neutral, fear, sadness) along with cognitive tasks, such as mathematical tasks or the Stroop Color Test (varies cognitive load through congruent and incongruent stimuli). They collected skin temperature, electrodermal activity, respiration, electrocardiogram, and photoplethysmogram data. The feature extraction was done on non-overlapping 45-second slices, which were normalized by subtracting the mean and dividing by the standard deviation of the training dataset. They then applied a set of machine learning classification algorithms. In contrast, we use data of a more complex real-world scenario, namely flight control, instead of an in-lab experimental setup. We assess a different aspect of subjective time, that is, the subjective perceived passage of time (time passage) instead of the relative estimation error based on verbal time estimation (duration). Our two main contributions are: (1) an automatic multi-modal subjective perceived passage of time perception estimation technique, using photoplethysmogram, electrodermal activity, and temperature data measured by of-the-shelf wearable sensors, with which, (2) we obtain high accuracies of up to 79 % on real-world physiological data obtained during flight controller helicopter training sequences. Our analysis includes an evaluation and discussion of eleven state-of-the-art machine learning classifiers for subjective perceived passage of time classification in combination with feature selection methods and feature importance analysis using the SHapley Additive exPlanations (SHAP) [16] framework. We show generalizability by validating our methods using leave-one-subject-out cross validation (LOSOCV).
@@ 1 H, GR1 H, EN2 H, GR2 H, ENvery slowslowmediumfastvery fastslowfast
Figure 2: Violin plots of the perceived passage of time labels of the different settings of the flight controller experiment (H: helicopter number, GR: Greek, EN: English; $n=12$). The left side of the violin plot (orange) displays the scaled label distribution from the questionnaire (scale is the left axis). The right side (blue) shows the label distribution after thresholding (dashed line) in the slow and fast class. The dots represent individual samples (overlapping). We observe that depending on the number of helicopters the majority of the controllers perceived the passage of time as slow when there is a one helicopter and fast when there are two helicopters.
For classification, we use the perceived passage of time estimation from the controllers’ questionnaires obtained at the end of the experiment. To allow for better comparability between questionnaire answers, we min-max-scale each controller’s answers from very slow to very fast, see Fig. 2, left half (orange). We sort the settings by speaking in Greek and one helicopter to speaking in English and two helicopters as we suspect an increase in mental load for speaking in a non-native language and an increased number of to-be-controlled objects. This sorting suggest a binary classification problem, namely, slow for one helicopter settings and fast for two helicopters settings. By thresholding the scaled labels at medium ($>3$ is labeled as fast), we obtain a binary classification problem with 26 sequences of class slow and 22 sequences of class fast, see Fig. 2, right half (blue).
@@ IV-4 Limitations
Considering the future application scenario of providing feedback to a ChronoPilot-device, we encounter several practical limitations for data collection. We require live-stream sensor data and a reliable communication protocol to reduce latency and to mitigate data loss. Also reducing data noise and removing artifacts need to work online. Depending on the use case, users may be required to be mobile (e.g., our industrial production scenario [3]). Motion artifacts may corrupt the physiological data, such as PPG. Hence, we may limit ourselves to physiological data that can be collected even if users are mobile and in motion. One possible PPG replacement is measuring electrocardiography (ECG), which directly measures the heart electrical activity. ECG allows for extracting the same features as PPG and there are methods to remove artifacts automatically. Using a thermopile as sensor also highly depends on the later use case, for example, considering again the industrial production scenario the user has to stand or walk around. One possible replacement could be measuring the skin and surrounding temperature. Skin temperature can often be measured together with EDA using the same device. Our results indicate that EDA is a strong descriptive feature and should definitely be used in further studies. Another interesting information source could be eye tracking, as it provides information about the user’s focus [24] and it could be integrated in a wearable ChronoPilot-device, such as VR/AR headsets.
@@ V Conclusion
We were able to show that it is feasible to estimate the perceived passage of time of air traffic controllers based on physiological data, such as PPG, EDA, and temperature measurements of a thermopile during helicopter training sequences of varying workload conditions with high accuracy up to 79 %. EDA seems to provide the most descriptive features based on SHAP analysis. This finding reinforces our assumption that we can influence the perceived passage of time by varying mental workload given that EDA is considered a biomarker for mental workload and stress. In terms of our objective of building a ChronoPilot-device, we have shown that we can use physiological data as feedback to close the control loop.
For our future work, larger datasets may improve the chances for generalizability. This can be done in two ways: increasing the number of participants or the amount of data per participant. In addition, larger datasets would allow us to explore deep learning methods. The literature on exploiting physiological data to automatically detect a person’s subjective time perception is scarce. To date, a closed-loop approach for guided manipulation of time perception has not yet been attempted. Our results reported here encourage us in our effort of creating an automatic closed-loop approach for modulating subjective time perception in real-world scenarios as envisioned in the ChronoPilot project. By exploiting the plasticity of subjective time perception, we can develop future technology that enhances productivity and well-being.